# references.bib

### v1

@@ 3.2 Implementation
Model. Our visual encoder is a MAE pre-trained ViT-B [he2022masked]. The input is rescaled to [0,1] and normalized with MAE’s averaged mean and std. Compared to the original ViT designed for 2D RGB images, our vision encoder only has three differences: (1) each scan is divided into 3D volumes instead of 2D patches; (2) the positional embedding encodes 3D spatial coordinates within each scan, along with an additional dimension to distinguish different scans; The cls token is minimally adapted to propagate information across layers that perform different hierarchies of attention.
@@
In our scenario, a curated dataset with only one imaging scan per study corresponds to the special case of M=1. For curated datasets, we employ slice attention with 4 evenly distributed scan attentions; for uncurated datasets, we use scan attention with 4 evenly distributed study attentions. We use PubMedBERT [pubmedbert] as our text encoder. Model configurations are provided in Appendix B.
@@
Preprocessing. We do not standardize the orientation or spacing. Instead, we consistently align the depth dimension with the through-plane axis of each scan, and then resize it to a fixed shape of (48,224,224). As this differs from prior practice [blankemeier2024merlin, hamamci2024developing, shui2025large], we provide a discussion in Appendix A.3. For brain MRI, we apply [0.5,99.5] percentile clipping to the intensity values. For head CT, we expand each scan with Hounsfield Unit (HU) values truncated to [0,120] for soft tissue, [-20,180] for contrast-enhanced tissue and blood vessels, and [-800,2000] for bone.
Pre-training. Our implementation builds upon OpenCLIP [cherti2023reproducible] and FLIP [li2023scaling]. We apply 25% patch dropout by default as regularization and acceleration. We do not apply additional augmentation or unmasked fine-tuning. Training 20 epochs on our uncurated datasets takes ~1 day with a batch size of 256 on 8 L40 GPUs. More details are provided in Appendix B.2.
@@ 4 Experiments
We apply HLIP to chest CT, brain MRI, and head CT. For chest CT and head CT, we follow prescribed evaluation protocols from previous work [hamamci2024developing, shui2025large, zhu20253d]. For brain MRI, we construct a new benchmark for zero-shot multi-class classification, using public datasets [dufumier2022openbhb, liu2023large, baid2021rsna, labella2023asnrmiccai, moawad2024brain, kazerooni2024brain, rudie2024university, link2024longitudinal]. To isolate the effectiveness of our hierarchical attention mechanism, we first apply HLIP to curated chest CT datasets [draelos2021machine, hamamci2024developing]. We then apply HLIP to our uncurated brain MRI and head CT datasets, highlighting the equivalent importance of scaling, modeling, and efficiency. Finally, we conduct comprehensive ablation and analysis to investigate our designs.
@@ 4.1 Chest CT
Implementation Details. Following CT-CLIP [hamamci2024developing] and fVLM [shui2025large], we apply HLIP to the CT-RATE [hamamci2024developing] training set, then perform internal validation on its test split and external validation on the full Rad-ChestCT dataset [draelos2021machine]. The implementation details differ from Section 3.2 as these datasets are curated. During preprocessing, each scan is standardized with a spacing of (3mm,1mm,1mm) [shui2025large]. The HU values are truncated to [-1150,350] [cao2024bootstrapping]. We apply a center crop of size (112,336,336) to construct mini-batch and a token size of (8,24,24). We use CXR-BERT [boecking2022making] as the text encoder. HLIP is then trained for 20 epochs without patch dropout, which can be completed within 6 hours using a batch size of 512 on 4 A40 GPUs. More details are provided in Appendix B.1.
Table 1: Results of multi-label zero-shot evaluation on the CT-RATE [hamamci2024developing] test split for internal validation and the Rad-ChestCT [draelos2021machine] for external validation. The best results are highlighted in bold.
@@
CT-CLIP [hamamci2024developing] 73.3 66.9 70.8 32.6 - - 63.3 59.9 64.7 34.1 - -
BIUD [cao2024bootstrapping] 71.3 68.1 71.6 33.8 67.3 62.9 60.6 65.2 33.7 59.6
Merlin [blankemeier2024merlin] 72.8 67.2 70.9 33.7 70.1 64.4 61.9 66.3 34.8 61.0
@@
Supervised by Qwen [bai2023qwen] Summarized Reports
fVLM [shui2025large] 77.8 71.8 75.1 37.9 72.8 68.0 64.7 68.8 37.4 64.6
@@
Results. Consistent with fVLM [shui2025large], we report the macro area under ROC curve (AUC); balanced accuracy (ACC); weighted F1-score (F1); recall; and precision for multi-label zero-shot classification in Table 1. When pre-trained with original reports, HLIP outperforms second-best models by 4.9% AUC on internal and 7.9% AUC on external validation. When pre-trained with reports summarized by large language models (e.g., Qwen [bai2023qwen], as used by fVLM), HLIP outperforms fVLM by 0.9% and 3.7% AUC on internal and external validation, respectively. The external generalizability of fVLM is largely limited by its reliance on anatomical segmentation, which, while facilitating the learning of fine-grained representations, introduces quality-related bias. HLIP, by contrast, demonstrates that a simple adaptation of the ViT backbone is sufficient to achieve superior performance.
During implementation, we observe that reports summarized by Qwen can improve performance on internal validation (i.e., +1.0% AUC), but do not necessarily lead to better performance on external validation (i.e., -0.6% AUC). This observation suggests that overly summarized supervision signals may actually constrain generalizability. This further prompted us to investigate the question: What additional factors could influence language-image pre-training in radiology? As this analysis lies beyond the main focus of the paper, we present the results in Appendix C.
4.2 Brain MRI
Pub-Brain-5. To the best of our knowledge, no publicly available benchmark exists for zero-shot transfer evaluation on brain MRI tasks. [^2] BrainMD [wang2024enhancing] is not publicly available due to an unexpected privacy policy. To this end, we construct a benchmark, named Pub-Brain-5, based on existing publicly available brain MRI datasets. Pub-Brain-5 comprises 18,343 studies drawn from Open-BHB [dufumier2022openbhb], the Stroke dataset [liu2023large], BraTS 2023 [baid2021rsna, labella2023asnrmiccai, moawad2024brain, kazerooni2024brain], NYU-Mets [link2024longitudinal], and UCSF-Mets [rudie2024university]. It spans five classes: healthy (3,984), acute stroke (2,871), glioma (1,614), meningioma (1,141), and metastasis (8,733). For each study, the number of scans ranges from 1 to 14. We further construct a subset of Pub-Brain-5, namely Pub-Brain-5-GT, which includes segmentation ground truth from the original datasets, enabling compatibility with medical foundation models that require 2D inputs [zhang2023biomedclip, nie2025conceptclip]. Pub-Brain-5-GT comprising 8,944 studies drawn from Open-BHB, the Stroke dataset, and BraTS 2023, covering the same five classes: healthy (3,984), acute stroke (2,372), glioma (1,350), meningioma (1,000), and metastasis (238). For both benchmarks, we evaluate three tasks: (i) binary disease detection (i.e., distinguishing diseased from healthy studies); (ii) three-class brain tumor classification; and (iii) five-class brain disease classification.
Baselines. Given that recent 3D models such as Merlin [blankemeier2024merlin] and M3D [bai2024m3d] do not include brain MRI in their training sets, we evaluate two 2D foundation models on our benchmark: BiomedCLIP [zhang2023biomedclip], pre-trained on 15 million figure-caption pairs from PubMed, and ConceptCLIP [nie2025conceptclip], pre-trained on 23 million such pairs. We use the prompt "This brain MRI shows: {disease}." to perform zero-shot inference [radford2021learning] with both models. Since both models require 2D inputs, we generate predictions for each study by applying average pooling and max pooling to the outputs across all slices. We observe that max pooling consistently yields positive predictions and is only suitable for binary classification; accordingly, we adopt average pooling for both models. Additionally, on Pub-Brain-5-GT, we assess an upper-bound scenario in which predictions are made only on lesion-containing slices. Note that this setting is not practical, as it requires manually annotation for the lesion locations.
Implementation Details. HLIP is trained on BrainMRI220K as described in Section 3.2, using the prompt "This MRI study shows: {disease}." during zero-shot transfer [radford2021learning].
Table 2: Results of zero-shot classification on Pub-Brain-5-GT and Pub-Brain-5. We report balanced accuracy (multi-class), with the best results highlighted in bold. "+annotation" indicates only using lesion-containing slices which are manually annotated by expert radiologists. These values are an upper-bound scenario for baselines, but not practical in real-world clinical settings.
Method Disease Detection Tumor Classification Disease Classification
Stroke Glioma Meningioma Metastasis
Pub-Brain-5-GT
BiomedCLIP [zhang2023biomedclip] 66.7 88.2 63.8 74.6 46.2 33.1
  +annotation 86.4 94.1 75.8 75.0 45.7 45.3
ConceptCLIP [nie2025conceptclip] 69.6 92.1 57.8 69.5 35.2 31.6
  +annotation 93.6 97.8 70.8 76.8 39.4 50.8
[] HLIP (ours) 95.0 89.2 79.6 73.4 54.8 61.3
Pub-Brain-5
BiomedCLIP [zhang2023biomedclip] 64.7 87.8 63.6 59.8 50.4 31.5
ConceptCLIP [nie2025conceptclip] 66.8 91.9 57.7 67.9 35.7 30.9
[] HLIP (ours) 91.5 89.2 79.2 78.1 63.3 63.9
Results. As two benchmarks are imbalanced, we report balanced accuracy for multi-class classification (mACC) in Table 2. On Pub-Brain-5-GT, we observe a performance boost for 2D foundation models when predicting on lesion-containing slices. Nevertheless, even when compared with the upper-bound scenario of baselines, HLIP still demonstrates superior performance in zero-shot brain tumor classification, i.e., outperforming BiomedCLIP [zhang2023biomedclip], by 8.6% mACC; and in brain disease classification, i.e., outperforming ConcepCLIP [nie2025conceptclip], by 10.5% mACC. When evaluated on the more comprehensive Pub-Brain-5 dataset, the performance gap for 2D foundation models widens due to the absence of lesion-containing annotations. Specifically, HLIP outperforms the second-best models by 12.9% mACC in brain tumor classification, and by 32.4% mACC in disease classification.
HLIP can occasionally lag behind BiomedCLIP [zhang2023biomedclip] and ConceptCLIP [nie2025conceptclip], particularly in the detection of brain glioma. However, we note that datasets like BraTS [baid2021rsna] are likely included in their pre-training corpora, as both models are trained on figure-caption pairs from PubMed. HLIP, on the other hand, does not encounter any instances from public datasets during training. The skull-stripping used in BraTS actually exacerbates the distribution shift. Moreover, HLIP makes predictions directly from the original radiology study, offering greater flexibility for real-world applications.
4.3 Head CT
Implementation Details. We apply HLIP to HeadCT240K, as described in Section 3.2. We then perform linear probing evaluation on two public benchmarks, RSNA [flanders2020construction] and CQ500 [chilamkurthy2018deep], exactly following FM-HeadCT [zhu20253d], which is trained on 361,633 head CT scans with DINOv2 [oquab2023dinov2].
(a) RSNA head CT hemorrhage dataset benchmarks a 5-way intracranial hemorrhage diagnosis task.
(b) CQ500 dataset benchmarks a 10-way task that includes hemorrhage diagnosis, bleed localization, midline shift, and mass effect detection.
Figure 3: Results of linear probing on the (a) RSNA [flanders2020construction] and (b) CQ500 [chilamkurthy2018deep] datasets. We report AUC, with our method shown in blue, FM-HeadCT [zhu20253d] in green, and Google CT [yang2024advancing] in red.
Results. Consistent with FM-HeadCT [zhu20253d], we report the area under the ROC curve (AUC) in Figure 3. On the RSNA dataset [flanders2020construction], HLIP outperforms the second-best model, FM-HeadCT [zhu20253d], on 4 out of 5 tasks, achieving an average AUC improvement of 1.4%. On the CQ500 dataset [chilamkurthy2018deep], HLIP outperforms both FM-HeadCT and Google-CT [yang2024advancing] on 7 out of 10 tasks, with average AUC improvements of 6.9% and 10.8%, respectively.
(a) Data Scale
(b) Radiology Study Modeling
(c) Computational Efficiency
Figure 4: Results of disease classification on Pub-Brain-5, highlighting the equal importance of (a) data scale, (b) radiology study modeling, and (c) computational efficiency. We report mACC.
(a) cls token strategy
(b) scan attn location
(c) # scan attn
@@ 4.3 Head CT
(e) cls token strategy
(f) study attn location
(g) # study attn
@@
Figure 5: Ablation study on the Rad-ChestCT [draelos2021machine] dataset (a)-(d) and the Pub-Brain-5 dataset (e)-(g). We report AUC and mACC for these two benchmarks, respectively.
4.4 Scaling, Modeling, and Efficiency
We acknowledge that one reason HLIP achieves superior performance (e.g., on Pub-Brain-5) is its training on a large-scale, domain-specific dataset. As shown in Figure 4(a), zero-shot disease classification on Pub-Brain-5 drops by 24.5% mACC when using only 10% of the training data. However, data scale is not the only contributing factor. Given an uncurated dataset, a straightforward approach is to randomly select one scan per study per step while training a 3D ViT. Modeling radiology studies in this way results in a 12.4% drop in mACC, even when trained on the full dataset with the same computational resources, as shown in Figure 4(b). Computational efficiency is equally important, as it determines how many samples can be contrasted within a batch. As shown in Figure 4(c), a small batch size (e.g., 64) results in a 6.6% drop in mACC. Therefore, HLIP bridges large-scale data, effective study modeling, and computational efficiency, enabling it to outperform current state-of-the-art foundation models [zhang2023biomedclip, nie2025conceptclip] by a substantial margin.
4.5 Ablations
Ablation on Rad-ChestCT. For the Rad-ChestCT dataset, our visual encoder is composed of slice attention layers with 4 evenly distributed scan attention layers. In Figure 5(a), we first ablate our strategy for propagating the cls token throughout the encoder, where none refers to applying average pooling at the end of the encoder, and learned refers to aggregate multiple cls tokens using attention pooling. We ablate the location of the scan attention layers in Figure 5(b), and the number of scan attention layers in Figure 5(c), where using 12 scan layers corresponds to the original ViT model—further highlighting the effectiveness of our hierarchical attention mechanism compared to ViT. In addition, we show that large batch size can achieve better performance in Figure 5(d).
Ablation on Pub-Brain-5. For the Pub-Brain-5 dataset, the encoder is composed of scan attention layers with 4 evenly distributed study attention layers. We ablate the cls token, the location and the number of study attention layers in Figure 5(e), Figure 5(f), and Figure 5(g), respectively. Figure 5(g) again highlights the effectiveness of our modeling compared to the original ViT.
4.6 Visualizations and Clinical Translation

### v2

@@ 3.2 Implementation
Model. Our visual encoder is a MAE pre-trained ViT-B [41]. The input is rescaled to [0,1] and normalized with MAE’s averaged mean and std. Compared to the original ViT designed for 2D RGB images, our vision encoder only has three differences: (1) each scan is divided into 3D volumes instead of 2D patches; (2) the positional embedding encodes 3D spatial coordinates within each scan, along with an additional dimension to distinguish different scans; The cls token is minimally adapted to propagate information across layers that perform different hierarchies of attention.
@@
For uncurated datasets, we use scan attention with 4 evenly distributed study attentions. A curated dataset with only one scan per study corresponds to the special case of M=1 in our scenario. For such datasets, we employ slice attention with 4 evenly distributed scan attentions; We use PubMedBERT [44] as our text encoder. Model configurations are provided in Appendix B.
@@
Preprocessing. We do not standardize the orientation or spacing. Instead, we consistently align the depth dimension with the through-plane axis of each scan, and then resize it to a fixed shape of (48,224,224). As this differs from prior practice [4, 3, 6], we provide a discussion in Appendix A.3. For brain MRI, we apply [0.5,99.5] percentile clipping to the intensity values. For head CT, we expand each scan into three separate scans, with Hounsfield Unit (HU) values truncated to [0,120] for soft tissue, [-20,180] for blood vessels, and [-800,2000] for bone.
Pre-training. Our implementation builds upon OpenCLIP [45] and FLIP [19]. We apply 25% patch dropout by default as regularization and acceleration. We do not apply additional augmentation or unmasked fine-tuning. Training 20 epochs on our uncurated datasets takes ˜1 day with a batch size of 256 on 8 L40 GPUs. More details are provided in Appendix B.2.
@@ 4 Experiments
We apply HLIP to brain MRI, head CT, and chest CT. For brain MRI, we construct a new benchmark for zero-shot classification, using public datasets [46, 23, 47, 48, 49, 50, 51, 52]. For head CT and chest CT, we follow evaluation protocols from previous work [3, 6, 20]. Experiments on our uncurated brain MRI and head CT datasets underscore the importance of effective modeling and demonstrating superior performance of HLIP over current foundation models [32, 14, 20, 53]. Experiments on curated chest CT datasets [24, 3] further isolate the effectiveness of the proposed hierarchical attention mechanism. Finally, we conduct comprehensive ablation and analysis to investigate our designs.
4.1 Brain MRI
Pub-Brain-5. To the best of our knowledge, no publicly available benchmark exists for zero-shot transfer evaluation on brain MRI tasks [^2] BrainMD [35] is not publicly available due to an unexpected privacy policy.. To this end, we construct a benchmark, named Pub-Brain-5, based on existing publicly available brain MRI datasets. Pub-Brain-5 comprises 18,343 studies drawn from Open-BHB [46], the Stroke dataset [23], BraTS 2023 [47, 48, 49, 50], NYU-Mets [52], and UCSF-Mets [51]. It spans five classes: healthy (3,984), acute stroke (2,871), glioma (1,614), meningioma (1,141), and metastasis (8,733). For each study, the number of scans ranges from 1 to 14. We also construct a subset of Pub-Brain-5, namely Pub-Brain-5-GT, which contains lesion-containing slice annotations derived from the original segmentation ground truth. Pub-Brain-5-GT comprising 8,944 studies drawn from Open-BHB, the Stroke dataset, and BraTS 2023, covering the same five classes: healthy (3,984), acute stroke (2,372), glioma (1,350), meningioma (1,000), and metastasis (238). We evaluate three zero-shot tasks: (i) binary anomaly detection (i.e., distinguishing diseased from healthy studies); (ii) three-way tumor classification; and (iii) five-way disease classification.
Baselines. We first evaluate the zero-shot transferability of two biomedical CLIP models on our benchmark: BiomedCLIP [32], pre-trained on 15 million figure-caption pairs from PubMed, and ConceptCLIP [14], pre-trained on 23 million such pairs. Since both models require 2D inputs, we generate study-level predictions by applying either average pooling or max pooling to the outputs across all slices. We observe that max pooling consistently produces positive predictions and is therefore only suitable for binary classification; accordingly, we adopt average pooling for both models. To enable a fairer comparison between these two models and HLIP, we assess an upper-bound scenario on Pub-Brain-5-GT, where predictions for the two models are made only on lesion-containing slices, while HLIP still takes the raw study as input. In addition, we evaluate PRIMA [54], which represents the vqvae + hierarchical transformer architecture family [3, 54], as well as a vanilla ViT that encodes the entire study.
Implementation Details. We use the prompt "This brain MRI shows: {disease}." to perform zero-shot inference [2] with biomedical CLIP models. HLIP and the vanilla ViT are trained on BrainMRI220K as described in Section 3.2, while PRIMA [54] is re-trained on the same dataset using the training recipe provided in the original paper. These three models employ the prompt "This MRI study shows: {disease.}" during zero-shot transfer.
Table 1: Results of zero-shot classification on Pub-Brain-5 and Pub-Brain-5-GT. We report balanced accuracy, with the best results highlighted in bold and the second-best result highlighted in underline. "+annotation" indicates only predicting on lesion-containing slices which are manually annotated.
Method Anomaly Detection
Tumor
Disease
Stroke Glioma Meningioma Metastasis mean
Pub-Brain-5
BiomedCLIP [32] 64.7 87.8 63.6 59.8 69.0 50.4 31.5
ConceptCLIP [14] 66.8 91.9 57.7 67.9 71.1 35.7 30.9
PRIMA [54] 78.8 89.3 70.8 64.7 75.9 42.8 31.6
ViT (our impl) 72.8 93.4 72.9 63.1 75.6 45.7 43.4
[] HLIP (ours) 91.5 89.2 79.2 78.1 84.5 63.3 63.9
Pub-Brain-5-GT
BiomedCLIP [32] 66.7 88.2 63.8 74.6 73.3 46.2 33.1
[]   +annotation 86.4 94.1 75.8 75.0 82.8 45.7 45.3
ConceptCLIP [14] 69.6 92.1 57.8 69.5 72.3 35.2 31.6
[]   +annotation 93.6 97.8 70.8 76.8 84.8 39.4 50.8
PRIMA [54] 61.2 81.0 87.7 53.4 70.8 45.9 31.4
ViT (our impl) 76.7 93.5 58.2 58.2 71.7 42.1 43.5
[] HLIP (ours) 95.0 89.2 79.6 73.4 84.3 54.8 61.3
Results. Given both benchmarks are imbalanced, we report balanced accuracy (ACC) in Table 1. On Pub-Brain-5, HLIP demonstrates superior performance over all baseline models, outperforming the second-best baseline by 20.5% ACC in disease classification. On Pub-Brain-5-GT, we observe a substantial performance boost for two biomedical CLIP models [32, 14] when predictions are restricted to lesion-containing slices, underscoring the fairness of our evaluation for these models even on uncurated studies. Notably, in disease classification, ConceptCLIP [14] achieves 50.8% ACC, outperforming PRIMA [54] and the vanilla ViT trained on our large-scale domain-specific datasets, indicating that merely scaling up data is not a sufficient solution for language-image pre-training in 3D medical imaging. On the other hand, HLIP achieves the new SOTA 61.3% ACC, attributable to the effectiveness of the proposed hierarchical attention mechanism.
4.2 Head CT
Implementation Details. We apply HLIP to HeadCT240K, as described in Section 3.2. For linear-probe evaluation, we benchmark HLIP on two public datasets, CQ500 [21] and RSNA [22], exactly following the protocol of FM-HeadCT [20]. For zero-shot evaluation, we evaluate HLIP on the full CQ500 and RSNA datasets and compare it with the vanilla ViT trained on the same dataset [^3] Recent CLIP models for 3D medical imaging, such as Merlin [4] and M3D [5], are excluded from our baselines, as they perform no better than random guessing in zero-shot evaluation on CQ500 and RSNA.. We employ the prompt "This CT study shows: {disease}." during zero-shot transfer.
Figure 3: Results of linear-probe and zero-shot evaluations on the CQ500 [21] and RSNA [22] datasets. We report AUC for each class, with blue represents our HLIP; green represents FM-HeadCT [20]; and red represents Google CT [53] in linear-probe evaluation and ViT in zero-shot evaluation.
Results. Consistent with FM-HeadCT [20], we report the area under the ROC curve (AUC) for each class in Figure 3. In linear-probe evaluation, HLIP demonstrates superior performance over FM-HeadCT [20], which is pre-trained on 361,663 studies using DINOv2 [55], and the Google-CT [53] foundation model. Specifically, on the CQ500 dataset [21], HLIP outperforms both FM-HeadCT and Google-CT with macro AUC improvements of 8.3% and 12.2%, respectively; on the RSNA dataset [22], HLIP achieves macro AUC improvements of 1.7% and 5.8% over the these baselines, respectively. In zero-shot evaluation, HLIP consistently outperforms the vanilla ViT, achieving macro AUC improvements of 10% on CQ500 and 2.4% on RSNA, further demonstrating the effectiveness of the proposed hierarchical attention mechanism.
@@ 4.3 Chest CT
Implementation Details. Following CT-CLIP [3] and fVLM [6], we apply HLIP to the CT-RATE [3] training set, then perform internal validation on its test split and external validation on the full Rad-ChestCT dataset [24]. The implementation details differ from those in Section 3.2. Following prior works [34, 6], we standardize each scan to a spacing of (3mm,1mm,1mm), truncate HU values to [-1150,350], construct mini-batches with a center crop of size (112,336,336), and apply a token size of (8,24,24). We also use CXR-BERT [7] as the text encoder. HLIP is trained for 20 epochs without patch dropout. Using a batch size of 512 on 4 A40 GPUs, training completes within 6 hours, significantly faster than prior works [3, 6], which require several days even with more advanced computational resources. More details are provided in Appendix B.1.
Table 2: Results of multi-label zero-shot evaluation on the CT-RATE [3] test split for internal validation and the Rad-ChestCT [24] for external validation. The best results are highlighted in bold.
@@
CT-CLIP [3] 73.3 66.9 70.8 32.6 - - 63.3 59.9 64.7 34.1 - -
BIUD [34] 71.3 68.1 71.6 33.8 67.3 62.9 60.6 65.2 33.7 59.6
Merlin [4] 72.8 67.2 70.9 33.7 70.1 64.4 61.9 66.3 34.8 61.0
@@
Supervised by Qwen [56] Summarized Reports
fVLM [6] 77.8 71.8 75.1 37.9 72.8 68.0 64.7 68.8 37.4 64.6
@@
Results. Consistent with fVLM [6], we report the area under ROC curve (AUC); balanced accuracy (ACC); weighted F1-score (F1); recall; and precision for multi-label zero-shot classification in Table 2. When pre-trained with original reports, HLIP outperforms second-best models by 4.9% AUC on internal and 7.9% AUC on external validation. When pre-trained with reports summarized by large language models (e.g., Qwen [56], as used by fVLM), HLIP outperforms fVLM by 0.9% and 3.7% AUC on internal and external validation, respectively. These results demonstrate that, compared with other components such as the VQ-VAE [57] tokenizer (used by CT-CLIP) and anatomical-guided fine-grained alignment (used by fVLM), the proposed hierarchical attention is a more effective and generalizable adaptation for language-image pre-training in 3D medical imaging.
(a) cls token prop. strategy
(b) study attn. location
(c) # study attn. layers
@@
(e) cls token prop. strategy
(f) scan attn. location
(g) # scan attn. layers
(h) batch size
@@
Figure 4: Ablation study on the Pub-Brain-5 (a)-(d) and Rad-ChestCT [24] dataset (e)-(h). We report balanced accuracy (ACC) on both datasets.
4.4 Ablation Study
In Figure 4, we present an ablation study on the key components of our visual encoder, including the cls token propagating strategy (Figure 4(a) and 4(e)); the location of dense attention (Figure 4(b) and 4(f)); and the number of dense attention layers (Figure 4(c) and 4(c)). We additionaly investigate the batch size in Figure 4(d) and 4(h), a factor overlooked in previous work [3, 6].
CLS token. When propagating the cls token across different attention levels, in addition to the default averaging setting, we also consider two alternative scenarios: (i) replacing the cls token with global average pooling at the end of the visual encoder (none); and (ii) applying weighted averaging using an attention pooling (learned). As shown in Figures 4(a) and 4(e), aggregating the cls token by averaging proves to be the most effective approach.
Location of dense attention layers. In Figure 4(b) and 4(f), we consider alternative locations of the dense attention layer: study attention for uncurated datasets, where each study contains multiple scans, and scan attention for curated datasets, where each study contains a single scan. We find that evenly distributed dense attention layers outperform other distributions, such as placing all at the front or all at the back.
Number of dense attention layers. In Figure 4(c) and 4(c), we investigate the number of dense attention layers, where 0 and 12 correspond to two implementations of vanilla ViT. We find that 4 dense attention layers are sufficient to capture global information, outperforming both 0 layers, which lack global information, and 12 layers, which lack the priors provided by slice or scan attention.
Batch size. Previous work [3, 6] used only small batch sizes (e.g., 48) during the pre-training. In Figure 4(d) and 4(h), we investigate this factor, which has been widely recognized as important in language-image pre-training, in the context of 3D medical imaging. We find that language-image pre-training in 3D medical imaging still benefits from larger batch sizes, underscoring the importance of memory-friendly designs such as our hierarchical attention mechanism.
4.5 Visualizations